twice the sum of a number and 7 is equal to three times the difference of the number and 2.
find the number

Answer by checkley77(12844)

2X+7=3(X-2)
2X+7=3X-6
2X-3X=-6-7
-X=-13
X=13 ANS.
PROOF:
2*13+7=3(13-2)
26+7=3*11
33=33